OpenStudy (anonymous):

kris can pain a room with a brush in 4 hours. gay can paint it with a roller in 2 hours. how long will it take the two to paint the room together

OpenStudy (anonymous):

Kris paints at a rate of 1 room / 4 hours = 0.25 rooms/hour Guy paints at a rate of 1 room / 2 hours = 0.5 rooms / hour Add the two rates together and you will obtain the combined rate 0.25 rooms/hour + 0.5 rooms/hour = 0.75 rooms / hour Switch this fraction around to get the number of hours per room at the combined rate.

OpenStudy (anonymous):

what if kris paints a room in 8 hours>?

hero (hero):

Hint: \frac{4 \times 2}{4 + 2} = x\] In general: \[\frac{M \times N}{M + N} = t\] Where: M = time it takes 1st person to do a job N = time it takes 2nd person to do a job t = time it takes while M and N are working together

OpenStudy (anonymous):

I like that. I think I will use it from now on.